# Value Profiles for Encoding Human Variation

### v1

@@ Abstract
Modelling human variation in rating tasks is crucial for enabling AI systems for personalization, pluralistic model alignment, and computational social science. We propose representing individuals using value profiles – natural language descriptions of underlying values compressed from in-context demonstrations – along with a steerable decoder model to estimate ratings conditioned on a value profile or other rater information. To measure the predictive information in rater representations, we introduce an information-theoretic methodology. We find that demonstrations contain the most information, followed by value profiles and then demographics. However, value profiles offer advantages in terms of scrutability, interpretability, and steerability due to their compressed natural language format. Value profiles effectively compress the useful information from demonstrations ($>$70% information preservation). Furthermore, clustering value profiles to identify similarly behaving individuals better explains rater variation than the most predictive demographic groupings. Going beyond test set performance, we show that the decoder models interpretably change ratings according to semantic profile differences, are well-calibrated, and can help explain instance-level disagreement by simulating an annotator population. These results demonstrate that value profiles offer novel, predictive ways to describe individual variation beyond demographics or group information.
@@ 1 Introduction
Machine learning systems are traditionally trained to approximate a single “ground truth” label, treating annotator disagreement as noise. However, many important tasks such as chat preferences, hate speech, and toxicity detection can have legitimate disagreement (Aroyo & Welty, 2015; Plank, 2022). Modelling this heterogeneity is important for pluralistic model alignment (Sorensen et al., 2024b), unbiased model safety, content moderation, personalization, and more.
We characterize three approaches to variation modelling: (1) Distributional Population Modelling, which directly models the distribution of labels for a given rater population (Zhang et al., 2024; Siththaranjan et al., 2024). This approach accounts for variance and valid disagreements between annotators but requires many raters labeling the same instances and doesn’t model which raters disagree or why. (2) Grouping by Characteristics such as demographics or annotation similarity. While grouping approaches can lead to higher agreement than the broader rater population, they still do not account for intra-group disagreement (Hwang et al., 2023; Prabhakaran et al., 2024), potentially leading to flattening variance or stereotyping. To capture intra-group variation, distributional learning is needed (Meister et al., 2024). Grouping by annotation similarity also requires significant overlap in labeled instances (Li et al., 2024; Vitsakis et al., 2024). (3) Individual Modelling. At the individual level (Gordon et al., 2022; Jiang et al., 2024), the target is a single ”correct” answer instead of a distribution111 At least, to the degree that people are self-consistent (Abercrombie et al., 2023)., allowing for standard supervised methods. Additionally, we can obtain group or population distributions through marginalization. Individual modeling also removes the requirement for raters to have any instance overlap. Because of these advantages, we argue for and focus on improving individual modeling in order to better model human variation However, this raises the question - how should we represent an individual?
In this work, we propose to model rater variation using individual, free-text value profiles – interpretable natural language descriptions of human values that explain observed rating variation (§2). In §3, we introduce a methodology to measure the information content of possible rater representations. We carry out a series of experiments to evaluate our value profile system and other rater representations (§4, 5). In §6, we introduce a rater clustering algorithm that uncovers better groupings than the most predictive demographics, while loosening the typical requirement of annotators labeling overlapping instances. In other experiments, we find that our value profile system is interpretable, well-calibrated, and helps explain rater disagreement (§7). We conclude by discussing related work, (§8), directions for future work (§9), and ethical advantages (and risks) of our approach (§10).
@@ 2.1 Rater Representations
Let =$\mathcal{R}=\{r_{1},r_{2},\ldots,r_{N_{R}}\}$ be the raters who we wish to model, $\mathcal{X}=\{x_{1},x_{2},\ldots,x_{N_{X}}\}$ be the space of instances, and $\mathcal{Y}=\{y_{1},y_{2},\ldots,y_{N_{Y}}\}$ be the space of potential responses/ratings. We would like to model $\mathcal{Y}\mid\mathcal{X},\mathcal{R}$. However, because we don’t have sufficient information to represent (or observe) the rater $r$, we compare different potential representations for $r$:
@@
  • $D$: Demographic information about $r$. $P(\mathcal{Y}\mid\mathcal{X},D(r))$ [^2] In the case that many demographics are provided, this is sometimes called a “persona” (Cheng et al., 2023). We also refer to this as “demographics (all)” or “intersectional demographics”. This is in contrast to trying to model an entire demographic group at a time. We compare against both settings.
@@ 2.2 Autoencoding Rater Values
Let $r_{i}$ be a particular rater $i$ drawn from the population of $n$ raters, $x_{j}$ be a particular instance $j$, and $y_{ij}$ be the rating that rater $i$ gave to instance $j$. Let $\mathcal{D}_{i}=\{y_{i1},y_{i2},\ldots,y_{iN_{i}}\}$ be the set of $N_{i}$ ratings we have for rater $i$. We can build a language model encoder $Q_{\phi}$ which estimates a value profile for each rater $r_{i}$ from a set of (fit) demonstrations drawn from $D_{i}$, with corresponding probability distribution $q_{\phi}:E_{n}\to V$. Similarly, a decoder $P_{\theta}$ can estimate the label probability distribution $P(\mathcal{Y}|\mathcal{X},V(\mathcal{R}))$approx $p_{\theta}:\mathcal{X},V\to\mathcal{Y}$. Given this, the entire autoencoder system can be evaluated by sampling a value profile from the encoder $v_{i}\sim q_{\phi}(E_{n}(r_{i}))$ and calculating the (cross entropy) loss on unseen examples.
@@
In practice, we initialize the encoder and decoder parameters $\phi$ and $\theta$ as prompted language models (prompts in Figs. A8/A9). For the experiments in the paper, we freeze the encoder parameters and optimize the decoder directly with supervised finetuning. We choose to do this 1) as prompted language models performed quite well at encoding, preserving $>70$% of usable information. (c.f. Eq. 3, Fig. 6), 2) in order to regularize the encoder to remain human understandable/interpretable, and 3) to preserve generalizability across datasets.
@@ 3 Estimating Usable Rater Information
Figure 3: An illustrative plot on fictional data for measuring $\mathcal{V}$-info.
@@
[…]
@@
$\bullet$ Training data $\mathcal{D}_{\text{train}}=\{(r_{i},x_{j},y_{ij})\}$ =$=\{(r_{i},x_{j},y_{ij}):(x_{j},y_{ij})\in R_{i}^{\text{test}},r_{i}\in\text{train raters }R^{\text{train}}\}$
$\bullet$ Test dataset $\mathcal{D}_{\text{test}}$ for held-out raters $R^{\text{test}}$, $R^{\text{train}}\cap R^{\text{test}}=\emptyset$
@@
$\bullet$ Some rater representation in natural language $g:R\to\text{NL}$
$d_{g}\leftarrow$ finetune $d$ on $\{(g(r_{i}),x_{j},y_{ij})|(r_{i},x_{j},y_{ij})\in\mathcal{D}_{\text{train}}\}$triangleright Train w/ rater information
$d_{\emptyset}\leftarrow$ finetune $d$ on $\{(\emptyset,x_{j},y_{ij})|(r_{i},x_{j},y_{ij})\in\mathcal{D}_{\text{train}}\}$triangleright Train w/out rater information
@@
$\hat{I}_{V}(g(R)\to\mathcal{Y}|\mathcal{X})\leftarrow H_{V}(\mathcal{Y}|\mathcal{X})-H_{V}(\mathcal{Y}|\mathcal{X},g(R))$triangleright Predictive information is drop in test loss when including rater information
@@
As Ethayarajh et al. (2022) show in a similar extension of $\mathcal{V}$-information, assuming we have an i.i.d. dataset of observations, we can get an unbiased estimate of this quantity for a computational family by training a model in each informational setting and comparing the held-out test losses to a trained model with no information. For more details, see Algorithm 1 (inspired by Ethayarajh et al. (2022)). To relate these concepts to an example loss plot, see Figure 3.
4 Experimental Methodology
Training details We split raters into 50/50 train/test splits and report results for training/test runs on five random splits. We draw $|\mathcal{D}_{i}^{\text{fit}}|\sim\mathcal{U}(\{2,\ldots,|\mathcal{D}_{i}|-2\})$ to ensure that we have variable-sized fit/eval splits with at least two instances each. We train the decoder (model: gemma2-9b-pt (Team et al., 2024b)) for a single epoch (important for maintaining calibration (Ji et al., 2021)). For encoders, we use gemma2-9b-it, gemma2-27b-it (Team et al., 2024b), and Gemini-1.5 Pro (Team et al., 2024a). See §11 for more details.
@@ 4 Experimental Methodology
Table 1: Dataset statistics including task information, number of multiple choice options (Ch.), demographic variables (Dem), unique instances (#I), unique raters (#R), and total ratings (#Rat). Datasets: OQA (Santurkar et al., 2023), HK (Kumar et al., 2021), DIC (Aroyo et al., 2023), VP (Sorensen et al., 2024a), HL (Tessler et al., 2024), and PR (Kirk et al., 2024b). Numbers may be smaller than in original datasets due to preprocessing/sampling (see §11). Results are noisier for datasets with $<$10k ratings due to underfit models/small test sets.
Datasets We utilize six datasets (Table 1) spanning tasks relevant to model alignment, content moderation, and computational social science. These datasets feature forced choice selection tasks and were selected to contain 1) some rater variation due to their subjective nature, 2) annotator IDs to link annotations from the same rater, and 3) ideally, some demographic information. Preprocessing information in §11.
@@ 5 Performance Across Rater Representation Settings
Detailed results for held-out test losses and accuracies across rater representations can be found in Figures 4 and A2 respectively.
Figure 4: Test losses across rater representation settings. Dashed line: label entropy $H(\mathcal{Y})$; no info: $\emptyset$; profile*: value profiles $V$ generated by gemma2-{9/27}b / Gemini-1.5-Pro; dem (all): $D$; N ex: $E_{N}$, up to $N$ examples from $D_{i}^{\textrm{fit}}$. ValuePrism does not have demographics, but does have a ground truth value profile. Each dot corresponds to a run with a differently seeded train/test split, with 95% CI reported. Significance is reported with a mixed-linear effects model for a difference with ”no info” (p $<$ .05/.01/.001). Generally, in-context examples are more performant than value profiles, which are more performant than demographics.
We note that error bars are much larger for 2 datasets, HL and PR. We believe that this is mainly because the datasets are smaller ($<$10k ratings), which means that 1) the trained model may be underfit and 2) there is a smaller sample size for each test split. We include results for all datasets for maximal inclusion, but focus our attention on the large datasets ($>$30k ratings: OQA, HK, DIC, VP) for which we can make higher confidence comparisons across settings.
Now, we compare decoder performance across rater representation settings (see Figs. 4, 5, A2). Our main findings are:
@@
[…]
@@
Figure 6: Info. preserved w.r.t. to using all examples. Results shown on the four large, low-variance datasets. Gemini value profiles preserve $>$70% of usable information.
@@
Value profiles effectively compressed the relevant information from in-context examples, preserving over 70% of the usable information for the four large datasets. This indicates that value profiles are a concise and efficient way to represent human variation.
Demographics have limited predictive power. Intersectional demographics generally offered a small and insignificant information boost, except for OpinionQA, where political affiliation was highly predictive. [^3] It makes sense that demographic variables offer a boost for OQA as political affiliation (included in demographics) can be highly predictive for a political opinion survey. Interestingly enough, however, the gains from demographic variables for other datasets were minimal. Additionally, value profiles contain more usable predictive information than demographics in all five datasets with demographics except OpinionQA (cf. Fig. 5)). This suggests that demographics alone may not be sufficient to capture the full spectrum of human variation.
@@ 6 Value Profile Clustering for Grouping Raters
We assign the train raters to clusters using Algorithm 2 (c.f. Figure 7), where each cluster corresponds to a single value profile description. [^4] While we focus on value profiles as cluster candidates, one could also use cluster candidates such as in-context examples or preset groups. We focus on value profiles due to their interpretability. We train a decoder to predict train rater annotations based on assigned cluster, and evaluate on held-out test raters. For all datasets, we use 100 randomly sampled value profiles as the cluster candidates. Results can be found in Figures 6 and 8 and the corresponding clusters can be found in Appendix H.
Figure 7: Clustering algorithm represented pictorially (also see Algorithm 2). 1) The decoder predicts label distributions for each instance and value profile combination; 2) calculate the loss for predicting each rater’s ”fit” ratings with each value profile; 3) find $C$ (# clusters) value profiles s.t. when each rater is assigned to a cluster, overall loss is minimized; 4) assign new raters to cluster with smallest loss on rater’s train ratings.

### v2

@@ Abstract
Modelling human variation in rating tasks is crucial for personalization, pluralistic model alignment, and computational social science. We propose representing individuals using natural language value profiles – descriptions of underlying values compressed from in-context demonstrations – along with a steerable decoder model that estimates individual ratings from a rater representation. To measure the predictive information in a rater representation, we introduce an information-theoretic methodology and find that demonstrations contain the most information, followed by value profiles, then demographics. However, value profiles effectively compress the useful information from demonstrations (>>70% information preservation) and offer advantages in terms of scrutability, interpretability, and steerability. Furthermore, clustering value profiles to identify similarly behaving individuals better explains rater variation than the most predictive demographic groupings. Going beyond test set performance, we show that the decoder predictions change in line with semantic profile differences, are well-calibrated, and can help explain instance-level disagreement by simulating an annotator population. These results demonstrate that value profiles offer novel, predictive ways to describe individual variation beyond demographics or group information.
Value Profiles for Encoding Human Variation
[…]
@@ 1 Introduction
Machine learning systems are traditionally trained to approximate a single “ground truth" label, treating annotator disagreement as noise. However, many important tasks such as chat preferences, hate speech, and toxicity detection can have legitimate disagreement (Aroyo and Welty, 2015; Plank, 2022). Modelling this heterogeneity is important for pluralistic model alignment (Sorensen et al., 2024b), unbiased model safety, content moderation, personalization, and more.
We characterize three approaches to variation modelling: (1) Distributional Population Modelling, which directly models the distribution of labels for a given rater population (Zhang et al., 2024; Siththaranjan et al., 2024). This approach accounts for variance and valid disagreements between annotators but requires many raters labeling the same instances and doesn’t model which raters disagree or why. (2) Grouping by Characteristics such as demographics or annotation similarity. While grouping approaches can lead to higher agreement than the broader rater population, they still do not account for intra-group disagreement (Hwang et al., 2023; Prabhakaran et al., 2024), potentially leading to flattening variance or stereotyping. To capture intra-group variation, distributional learning is needed (Meister et al., 2024). Grouping by annotation similarity also requires significant overlap in labeled instances (Li et al., 2024). (3) Individual Modelling. At the individual level (Gordon et al., 2022; Jiang et al., 2024), the target is a single "correct" answer instead of a distribution, [^1] At least, to the degree that people are self-consistent (Abercrombie et al., 2023). allowing for standard supervised methods. Additionally, we can obtain group or population distributions through marginalization. Individual modeling also removes the requirement for raters to have any instance overlap. Because of these advantages, we argue for and focus on improving individual modeling in order to better model human variation (for more, see App. B, Fig. A1). However, this raises the question - how should we represent an individual?
@@
In this work, we propose to model rater variation using individual, free-text value profiles – interpretable natural language descriptions of human values that explain observed rating variation (§2). In §3, we introduce a methodology to measure the information content of possible rater representations. We carry out a series of experiments to evaluate our value profile system and other rater representations (§4, 5). In §6, we introduce a rater clustering algorithm that uncovers better groupings than the most predictive demographics, while loosening the typical requirement of annotators labeling overlapping instances. In other experiments, we find that our value profile system is interpretable, well-calibrated, and helps explain rater disagreement (§7). We conclude by discussing related work (§8), directions for future work (§9), and ethical advantages (and risks) of our approach (§11).
@@ 2.1 Rater Representations
Let $\mathcal{R}=\{r_{1},r_{2},\ldots,r_{N_{R}}\}$ be the raters who we wish to model, $\mathcal{X}=\{x_{1},x_{2},\ldots,x_{N_{X}}\}$ be the space of instances, and $\mathcal{Y}=\{y_{1},y_{2},\ldots,y_{N_{Y}}\}$ be the space of potential responses/ratings. We would like to model $\mathcal{Y}\mid\mathcal{X},\mathcal{R}$. However, because we don’t have sufficient information to represent (or observe) the rater $r$, we compare different potential representations for $r$:
@@
  • $D$: Demographic information about $r$. $P(\mathcal{Y}\mid\mathcal{X},D(r))$ [^2] In the case that many demographics are provided, this is sometimes called a “persona” (Cheng et al., 2023). We also refer to this as “demographics (all)” or “intersectional demographics”. This is in contrast to trying to model an entire demographic group at a time.
@@ 2.2 Autoencoding Rater Values
Let $r_{i}$ be a particular rater $i$ drawn from the population of $n$ raters, $x_{j}$ be a particular instance $j$, and $y_{ij}$ be the rating that rater $i$ gave to instance $j$. Let $\mathcal{D}_{i}=\{y_{i1},y_{i2},\ldots,y_{iN_{i}}\}$ be the set of $N_{i}$ ratings we have for rater $i$. We can build a language model encoder $Q_{\phi}$ which estimates a value profile for each rater $r_{i}$ from a set of (fit) demonstrations drawn from $D_{i}$, with corresponding probability distribution $q_{\phi}:E_{n}\to V$. Similarly, a decoder $P_{\theta}$ can estimate the label probability distribution $P(\mathcal{Y}|\mathcal{X},V(\mathcal{R}))$ $\approx$ $p_{\theta}:\mathcal{X},V\to\mathcal{Y}$. Given this, the entire autoencoder system can be evaluated by sampling a value profile from the encoder $v_{i}\sim q_{\phi}(E_{n}(r_{i}))$ and calculating the (cross-entropy) loss on unseen examples.
@@
In practice, we initialize the encoder and decoder parameters $\phi$ and $\theta$ as prompted language models (prompts in Figs. A9/A10). For the experiments in the paper, we freeze the encoder parameters and optimize the decoder directly with supervised finetuning. We choose to do this 1) as prompted language models performed quite well at encoding, preserving >70>70% of usable information. (cf. Eq. 3, Fig. 5), 2) in order to regularize the encoder to remain human understandable/interpretable, and 3) to preserve generalizability across datasets.
@@ 3 Estimating Usable Rater Information
$H_{\mathcal{V}}(B\mid A,C)=\inf_{f\in\mathcal{V}}\mathbb{E}_{a,b,c\sim A,B,C}[-\log f[a,c](b)]$ (1)
$I_{\mathcal{V}}(A\to B\mid C)=H_{\mathcal{V}}(B\mid\emptyset,C)-H_{\mathcal{V}}(B\mid A,C)$ (2)
As Ethayarajh et al. (2022) show in a similar extension of $\mathcal{V}$-information, assuming we have an i.i.d. dataset of observations, we can get an unbiased estimate of this quantity for a computational family by training a model in each informational setting and comparing the held-out test losses to a trained model with no information. For more details, see Algorithm 1 (inspired by Ethayarajh et al. 2022). To contextualize the algorithm with an example loss plot, see Figure A2.
@@
$\bullet$ Training data $\mathcal{D}_{\text{train}}=\{(r_{i},x_{j},y_{ij})\}$ $\{(r_{i},x_{j},y_{ij}):(x_{j},y_{ij})\in R_{i}^{\text{test}},r_{i}\in\text{train raters }R^{\text{train}}\}$
$\bullet$ Test data $\mathcal{D}_{\text{test}}$ for held-out raters $R^{\text{test}}$, $R^{\text{train}}\cap R^{\text{test}}=\emptyset$
@@
$\bullet$ Natural language rater representation $g:R\to\text{NL}$
$d_{g}\leftarrow$ finetune $d$ on $\{(g(r_{i}),x_{j},y_{ij})|(r_{i},x_{j},y_{ij})\in\mathcal{D}_{\text{train}}\}$ $\triangleright$ Train w/ rater information
$d_{\emptyset}\leftarrow$ finetune $d$ on $\{(\emptyset,x_{j},y_{ij})|(r_{i},x_{j},y_{ij})\in\mathcal{D}_{\text{train}}\}$ $\triangleright$ Train w/out rater information
@@
$\hat{I}_{V}(g(R)\to\mathcal{Y}|\mathcal{X})\leftarrow H_{V}(\mathcal{Y}|\mathcal{X})-H_{V}(\mathcal{Y}|\mathcal{X},g(R))$ $\triangleright$ Predictive information is drop in test loss when including rater information
@@
Table 1: Dataset statistics including task information, number of multiple choice options (Ch.), demographic variables (Dem), unique instances (#I), unique raters (#R), and total ratings (#Rat). Datasets: OQA (Santurkar et al., 2023), HK (Kumar et al., 2021), DIC (Aroyo et al., 2023), VP (Sorensen et al., 2024a), HL (Tessler et al., 2024), and PR (Kirk et al., 2024b). Numbers may be smaller than in original datasets due to preprocessing/sampling (see §A). Results are noisier for datasets with <<10k ratings due to underfit models/small test sets.
4 Experimental Methodology
Training details We split raters into 50/50 train/test splits and report results for training/test runs on five random splits. We draw $|\mathcal{D}_{i}^{\text{fit}}|\sim\mathcal{U}(\{2,\ldots,|\mathcal{D}_{i}|-2\})$ to ensure that we have variable-sized fit/eval splits with at least two instances each. We train the decoder (gemma2-9b-pt, Gemma Team et al. 2024) for a single epoch (important for maintaining calibration, Ji et al. 2021). For encoders, we use gemma2-9b-it, gemma2-27b-it (Gemma Team et al., 2024), and Gemini-1.5 Pro (Team et al., 2024). See App. A for details.
Datasets We utilize six datasets intended for research (Table 1) spanning tasks relevant to model alignment, content moderation, and computational social science. These datasets feature forced choice selection tasks and were selected to contain 1) some rater variation due to their subjective nature, 2) annotator IDs to link annotations from the same rater, and 3) ideally, some demographic information. Preprocessing information in §A.
Figure 3: Test losses across rater representation settings. Dashed line: label entropy $H(\mathcal{Y})$; no info: $\emptyset$; profile*: value profiles $V$ generated by gemma2-{9/27}b / Gemini-1.5-Pro; dem (all): $D$; N ex: $E_{N}$, up to NN examples from $D_{i}^{\textrm{fit}}$. ValuePrism does not have demographics, but does have a ground truth value profile. Each dot corresponds to a run with a differently seeded train/test split, with 95% CI reported. Generally, in-context examples are more performant than value profiles, which are more performant than demographics.
@@ 5 Performance Across Rater Representation Settings
Detailed results for held-out test losses across rater representations can be found in Figure 3. Accuracies can be found in App. A3, but results mirror the loss results, which we will focus on. Detailed results for held-out test losses and accuracies across rater representations can be found in Figures 3 and A3 respectively.
We note that error bars are much larger for 2 datasets, HL and PR. We believe that this is mainly because the datasets are smaller (<<10k ratings), which means that 1) the trained model may be underfit and 2) there is a smaller sample size for each test split. We include results for all datasets for maximal inclusion, but focus our attention on the large datasets (>>30k ratings: OQA, HK, DIC, VP) for which we can make higher confidence comparisons across settings.
Now, we compare decoder performance across rater representation settings (see Figs. 3, 4, A3). Our main findings are:
@@
In-context examples improve predictions. Across all four large datasets, providing the decoder with in-context examples of the rater’s previous annotations significantly improved the prediction of their ratings on held-out test data in both accuracy and test loss ($p<.001$). We observe a similar, but less significant, drop in loss/increase in accuracy on the two small datasets. This shows that rater demonstrations offer useful information for disentangling human variation.
@@
Value profiles effectively compressed the relevant information from in-context examples, preserving >>70% of the usable information for the four large datasets. This indicates that value profiles are an efficient way to represent human variation.
Demographics have limited predictive power. Intersectional demographics generally offered a small and insignificant information boost, except for OpinionQA, where political affiliation was highly predictive. [^3] It makes sense that demographic variables offer a boost for OQA as political affiliation (included in demographics) can be highly predictive for a political opinion survey. Interestingly enough, however, the gains from demographic variables for other datasets were minimal. Additionally, value profiles contain more usable predictive information than demographics in all five datasets with demographics except OpinionQA (cf. Fig. 4). This suggests that demographics alone may not be sufficient to capture the full spectrum of human variation.
@@
Figure 5: Info. preserved w.r.t. to using all examples. Results shown on the four large, low-variance datasets. Gemini profiles preserve >>70% of usable information.
@@ 6 Value Profile Clustering for Grouping Raters
Figure 6: Clustering algorithm represented pictorially (also see Algorithm 2). 1) The decoder predicts label distributions for each instance and value profile combination; 2) calculate the loss for predicting each rater’s "fit" ratings with each value profile; 3) find $C$ (# clusters) value profiles s.t. when each rater is assigned to a cluster, overall loss is minimized; 4) assign new raters to cluster with smallest loss on rater’s train ratings.
@@
We assign the train raters to clusters using Algorithm 2 (cf. Figure 6), where each cluster corresponds to a single value profile description. [^4] While we focus on value profiles as cluster candidates, one could also use cluster candidates such as in-context examples or preset groups. We focus on value profiles due to their interpretability. We train a decoder to predict train rater annotations based on assigned cluster, and evaluate on held-out test raters. For all datasets, we use 100 randomly sampled value profiles as the cluster candidates. Results can be found in Figs. 5/7 and the corresponding clusters can be found in Appendix H.